OpenStudy (anonymous):

The Remainder Theorem states that: Select one: a. if a polynomial f(x) is added by x – k, then the remainder is r = f(k). b. if a polynomial f(x) is multiplied by x – k, then the remainder is r = f(k). c. if a polynomial f(x) is divided by x – k, then the remainder is r = f(k). d. none of the above

pooja195 (pooja195):

"The Remainder Theorem When we divide a polynomial f(x) by x-c we get: f(x) = (x-c)·q(x) + r(x) But r(x) is simply the constant r (remember? when we divide by (x-c) the remainder is a constant) .... so we get this: f(x) = (x-c)·q(x) + r Now see what happens when we have x equal to c: f(c) = (c-c)·q(c) + r f(c) = (0)·q(c) + r f(c) = r" https://www.mathsisfun.com/algebra/polynomials-remainder-factor.html With this said what do you think the answer is? Sorry if it seems confusing..

OpenStudy (anonymous):

multiplied?

OpenStudy (jadzia):

Take note of the word "\(\sf \color{red}{Remainder}\)" and try to read what pooja said about the "\(\sf \color{red}{Remainder}\)" Theorem. What should you get to get a remainder? do you add? do you multiply? do you subtract? or do you divide?

OpenStudy (anonymous):

It states that the remainder of the division of a polynomial by a linear polynomial is equal to. In particular, is a divisor of if and only if. so division?

OpenStudy (anonymous):

yes